# Performance studies of the CE-65v2 MAPS prototype structure

### v3

@@ Abstract
With the next upgrade of the ALICE inner tracking system (ITS3) as its primary focus, a set of small MAPS test structures have been developed in the 65 nm TPSCo CMOS process. The CE-65 focuses on the characterisation of the analogue charge collection properties of this technology. The latest iteration, the CE-65v2, was produced in different processes (standard, with a low-dose n-type blanket, and blanket with gap between pixels), pixel pitches (15, 18, ), and pixel arrangements (square or staggered). The comparatively large pixel array size of pixels in CE-65v2 allows, among other benefits, to study the uniformity of the pixel response.
@@ 1 Introduction
Radiation tolerance | Few per year | |
@@ 5 In-pixel efficiency and mean absolute deviation studies
The in-pixel measurement of the mean absolute deviation between the reconstructed telescope and DUT cluster positions () uses the window method. Fig. 3 shows the mean absolute deviation depending on the position of the hit inside the pixel. In the STD process, the distribution is more uniform and does not exceed anywhere in the pixel. For the GAP sensor, the performance strongly deteriorates at the edges and corners, which leads to the overall worse spatial resolution observed before. Using the window method, the full power of the STD process is showcased: Charge sharing from the seed to the neighbouring pixels strongly helps to reconstruct the hit position and reach resolutions much below . ion*Conclsions and Outlook
@@
The challenges of achieving resolution in a l large digital read-out chip efficient at reasonable noise levels and in a radiation environment differ between the two processes. In , an intricate analogue-to-digital conversion scheme is necessary to fully benefit from the charge sharing information while keeping the readout data rate minimal. For the GAP process, the challenge is to integrate all the necessary logic into pixels of small pitch required to achieve resolution.

### v4

@@ Abstract
With the next upgrade of the ALICE inner tracking system (ITS3) as its primary focus, a set of small MAPS test structures have been developed in the 65 nm TPSCo CMOS process. The CE-65 focuses on the characterisation of the analogue charge collection properties of this technology. The latest iteration, the CE-65v2, was produced in different processes (standard, with a low-dose n-type blanket, and blanket with gap between pixels), pixel pitches (15, 18, ), and pixel arrangements (square or staggered). The comparatively large pixel array size of pixels in CE-65v2 allows the uniformity of the pixel response to be studied, among other benefits.
@@ 1 Introduction
Radiation tolerance | Several per year | |
@@ 5 In-pixel efficiency and mean absolute deviation studies
The in-pixel measurement of the mean absolute deviation between the reconstructed telescope and DUT cluster positions () uses the window method. Fig. 3 shows the mean absolute deviation depending on the position of the hit inside the pixel. In the STD process, the distribution is more uniform and does not exceed anywhere in the pixel. For the GAP sensor, the performance strongly deteriorates at the edges and corners, which leads to the overall worse spatial resolution observed before. Using the window method, the full power of the STD process is showcased: Charge sharing from the seed to the neighbouring pixels strongly helps to reconstruct the hit position and reach resolutions much below .
6 Conclusions and Outlook
@@
The challenges of achieving resolution in a l large digital read-out chip efficient at reasonable noise levels and in a radiation environment differ between the two processes. In STD, an intricate analogue-to-digital conversion scheme is necessary to fully benefit from the charge sharing information while keeping the readout data rate minimal. For the GAP process, the challenge is to integrate all the necessary logic into pixels of small pitch required to achieve resolution.